Solve the equation 2x3 + x2 - 13x + 6=0.
One of the roots lies between 1 and 3.

Answer by lwsshak3(11628)

Solve the equation 2x3 + x2 - 13x + 6=0.
One of the roots lies between 1 and 3.
..
If you are not allowed to use the graphing calculator, the best thing to do is try x=2 as one of the roots. Divide x-2 into the function, 2x3 + x2 - 13x + 6, by long division or synthetic division (sorry, this format does not allow me to show you how to do synthetic division, but perhaps you already know how.) You will then get a quotient, 2x^2+5x-3. You now have:
(x-2)(2x^2+5x-3)=0
(x-2)(2x+6)(x-1/2)=0
ans: function has three real roots:
x=2
x=1/2
x=-3